# HyperPg - Prototypical Gaussians on the Hypersphere for Interpretable Deep Learning

### v1

@@ 7 Conclusion
This work introduces HyperPg, a new prototype representation learning a probability distribution on the surface of a hypersphere in latent space. HyperPg adapt to the variance of clusters in latent space and improve training time and accuracy compared to other prototype formulations. HyperPgNet leverages HyperPg to learn human defined concept prototypes, instead of black-box optimized prototypes. The combination of probabilistic prototypes on the hypersphere and concept aligned prototypes allows HyperPgNet to outperform other prototype learning approaches. One limitation HyperPgNet faces are increased computational requirements due to the inclusion of concept annotations and $L_{\mathrm{RRC}}$ during training. However, this is offset by an acceleration of the training process regarding epochs and number of samples. Coupled with the increased transparency and interpretability of the model, this makes HyperPgNet a strong contender for scenarios with few training samples and high safety requirements, like medical applications or human-robot interaction.
Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 448648559
@@ A.1 Cauchy Distribution
Fig. 7 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$. The Gaussian distributions PDF quicly approaches zero and stays near constant. This could potentially cause vanishing gradient issues during training. The heavier tails of the Cauchy distribution ensure that for virtually the entire value range of the cosine similarity, gradients could be propagated back through the model. However, experiments on CUB-200-2011 showed no significant performance difference between using HyperPg with the Gaussian or Cauchy distribution.
@@ A.2 Truncated Distributions
$\propto\exp\left(\kappa{\bm{v}}^{\top}\bm{{\alpha}}\right).$ (24)
Eq. 23 also shows the relationship to the HyperPg similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$.
@@ A.5 Mixture Models
Appendix B Ablations
B.1 Backbone Performance
HyperPgNets performance is not dependent on a CNN or Transformer based backbone. Fig. 10 shows the average test accuracy curves for HyperPgNet with concept aligned prototypes on CUB-200-2011. The maximum differnece in test accuracy between both architectures is 2%.
Figure 10: CUB-200-2011 Test Accuracies with different Backbones. The maximum difference in performance is 2%.
B.2 Number of prototypes
The number of prototypes per class or concept is an important hyper-parameter for prototype learning models like HyperPgNet. The choice of the number of prototypes controls, how many clusters the model should fit in latent space. For example, ProtoPNet requires 10 prototypes per class or 2000 prototypes in total for the CUB-200-2011 dataset (Chen et al., 2019). ProtoPool is able to merge similar prototypical parts, requiring only 202 prototypes for its best performance (Rymarczyk et al., 2022).
The change from class specific to concept aligned prototypes improves overall test performance, and the added information during training leads to fewer training epochs. Concept Bottleneck Models (CBM) are able to outperform the other models with only 30 learned concepts, or 1 prototype per concept. This indicates that the concept annotations on CUB-200-2011 do not produce multi-modal distributions in latent space, which would require more prototypes per concept. This behavior is confirmed in ablation studies on the number of prototypes for HyperPgNet. Fig. 11 presents the test accuracies for HyperPgNet on full CUB-200-2011 images after 60 training epochs with different numbers of prototypes per concept.
Figure 11: Ablation on number of prototypes in HyperPgNet on the CUB-200-2011 dataset. CUB-200-2011’s concept annotations have sufficient granularity for modeling with one prototype per concept. This is in line with CBM’s perfomance, which also learns only one prototype per concept.
@@ Appendix C Extended Interpretability Analysis
C.1 Right for the Right Concept
Figure 12: Concept Prototypes from HyperPgNet
Fig. 12 shows more example prototypes for HyperPgNet. Without $L_{\mathrm{RRC}}$, the prototype do not always correspond to the defined prototype. For the “Orange Body” concept, the prototype also activates on confounding factors in the background. With $L_{\mathrm{RRC}}$ during training, the model learns to ignore these factors.
With the car dataset, the effect is even more pronounced. First analyses on the Stanford Cars experiments indicate, that most models focus on the front bumper below the headlights. This part of the car seems to be highly predictive for the class label. Because of this, the HyperPgNet, even with context annotations but without $L_{\mathrm{RRC}}$ will focus more on these areas of the image. For example the “bonnet” prototype does not focus on the car bonnet at all. Only when trained with $L_{\mathrm{RRC}}$, the prototype actually focuses on the human defined concept.
C.2 Latent Space Structure
Prototype learning is based on learning structures in the latent space. HyperPg specifically learns Gaussian distributions on the surface of a hypersphere in a high-dimensional latent space. Dimensionality reduction techniques like UMAP (McInnes et al., 2018) aim to preserve global and local structures from a high dimensional space when projecting into a low dimensional one. UMAP supports multiple distance metrices, enabling the projection onto the surface of a 3D sphere.
@@ C.2 Latent Space Structure
Fig. 14 illustrates one UMAP projection of HyperPg concept-aligend prototypes trained on CUB-200-2011. In comparison to Stanford Cars, the latent space appears less structured. This could explain the higher difficulty associated with this dataset.
@@
Figure 13: UMAP projection of HyperPg concept prototypes learned on Stanford Cars. Each dot represents one prototype, colored by the concept it represents.

### v2

@@ 7 Conclusion
This work introduces HyperPg, a new prototype representation learning a probability distribution on the surface of a hypersphere in latent space. HyperPg prototypes adapt to the variance of clusters in latent space and improve training time and accuracy compared to other prototype formulations. HyperPgNet leverages HyperPg to learn human-defined concept prototypes instead of black-box optimized prototypes. The combination of probabilistic prototypes on the hypersphere and concept-aligned prototypes allows HyperPgNet to outperform other prototype learning approaches with regards to accuracy and interpretability. One limitation HyperPgNet faces are slightly higher computational requirements due to the inclusion of concept annotations and $L_{\mathrm{RRC}}$ during training. However, this is offset by faster convergence. Coupled with the increased transparency and interpretability of the model, this makes HyperPgNet a strong contender for scenarios with higher requirements for model trust and safety, like medical applications or human-robot interaction.
Reproducibility Statement
For reproducing the experiments, please refer to the following sections: Sec. 4 describes the required loss functions and the overall network architecture. Subsection 3.3 details the activation function of the newly introduced HyperPg prototypes. Appendix B provides detailed information on the experiment implementations, including training hyperparameters, processing of the datasets and model implementation details. Furthermore, we will publish the code on GitHub after the double-blind review period.
@@ A.1 Cauchy Distribution
Fig. 9 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$. The Gaussian distributions PDF quickly approaches zero and stays near constant. This could potentially cause vanishing gradient issues during training. The heavier tails of the Cauchy distribution ensure that for virtually the entire value range of the cosine similarity, gradients could be propagated back through the model. However, experiments on CUB-200-2011 showed no significant performance difference between using HyperPg with the Gaussian or Cauchy distribution.
@@ A.2 Truncated Distributions
$=\exp(\kappa{\bm{v}}^{\top}\bm{{\alpha}}-\kappa)$ (24)
$=\exp(\kappa)^{-1}\exp(\kappa{\bm{v}}^{\top}\bm{{\alpha}})$ (25)
$\sim\exp\left(\kappa{\bm{v}}^{\top}\bm{{\alpha}}\right).$ (26)
Eq. 23 also shows the relationship to the HyperPg similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$. Fig. 11 presents a simulation of the vMF distribution on a 3D sphere. While both the vMF distribution and HyperPg activation can produce a spherical, gaussian-like activation pattern on the surface of a hypersphere, the vMF distribution cannot produce the ring pattern shown in Fig. 2. The ring pattern produced by adapting HyperPg’s mean similarity $\mu$ could be approximated by a mixture of vMF distributions.
Figure 11: Changing the concentration parameter $\kappa$ is akin to changing HyperPg’s std $\sigma$.
@@ A.5 Mixture Models
Appendix B Implementation Details
B.1 Data Preprocessing
In contrast to prior work (e.g. Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch. The input images were first resized to a resolution of $224\times 224$ without cropping to bounding box annotations. The augmentations consisted of RandomPerspective, RandomHorizontalFlip and RandomAffine.
B.2 Hyperparameters
The prototypical part networks ProtoPNet and HyperPgNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 256$. The models were trained with a minibatch size of 96 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
B.3 Compute Resources
All experiments were performed on a workstation with a single NVIDIA RTX 3090 GPU (24 GB VRAM) per model.
B.4 Model Implementation
The models were implement with different feature encoding backbones. ConvNeXt-tiny (Liu et al., 2022) and ResNet50 (He et al., 2016) with pretrained weights on ImageNet provide CNN based backbones. As a transformer based backbone MiT-B4 (Xie et al., 2021) as provided by HuggingFace is used. MiT-B4 is pretrained on ImageNet and finetuned on ADE20k, Cityscapes and COCO-stuff.
The models are implemented as follows:
  • The black box baseline model uses the pretrained feature encoder as backbone followed by the classification head: a single linear output layer and softmax activation.
  • Concept Bottleneck Models (Koh et al., 2020) integrate the concept bottleneck, a linear layer with one neuron per concept, between the backbone and classification head.
  • ProtoPNet (Chen et al., 2019) use a $L_{2}$ based prototype layer between backbone and classification head.
  • ProtoPNet + HyperPg use a HyperPg based prototype module between backbone and classification head.
  • HyperPgNet $-{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head. The prototypes are trained with concept alignment, but without the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
  • HyperPgNet $+{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head. The prototypes are trained with concept alignment and including the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
@@ Appendix C Extended Interpretability Analysis
C.1 Latent Space Structure
Prototype learning is based on learning structures in the latent space. HyperPg specifically learns Gaussian distributions on the surface of a hypersphere in a high-dimensional latent space. Dimensionality reduction techniques like UMAP (McInnes et al., 2018) aim to preserve global and local structures from a high dimensional space when projecting into a low dimensional one. UMAP supports multiple distance metrics, including hyperspheric manifold distances, thus retaining some of the high dimensional structure when projecting onto a 3D sphere.
@@ C.1 Latent Space Structure
Fig. 14 illustrates one UMAP projection of HyperPg concept-aligned prototypes trained on CUB-200-2011. In comparison to Stanford Cars, the latent space appears less structured. This could explain the higher difficulty associated with this dataset.
@@
Figure 13: UMAP projection of HyperPg concept prototypes learned on Stanford Cars. Each dot represents one prototype with the color indicating a concept.